OpenStudy (anonymous):

Which property justifies the following equation? 7(6 + 0 + 5) = 7(6 + 5) A. distributive B. commutative C. associative D. identity

OpenStudy (anonymous):

help

OpenStudy (anonymous):

the last one property called as identity according to which (0+k)=k

OpenStudy (anonymous):

?

OpenStudy (texaschic101):

it is D. Identity property states that any number + 0 = the original number.

OpenStudy (anonymous):

for any constant k k+0=k is called identity therefor 7(6+0+5)=7(6+5) also called identity property hence d